# Topology of Discrete Quantum Feedback Control

### v2

@@ III Topological equivalence of quantum channels
In this section, we introduce topology of quantum channels (CPTP maps) to investigate topological phases induced by feedback control.
@@ III.1 Setup
be a set of lattice sites, where $\vec{a}_{\lambda}\ (\lambda=1,\cdots,d)$ denote the primitive lattice vectors. Each lattice site $\vec{R}_{\vec{i}}$ is specified by a set of integers $\vec{i}=(i_{1},\cdots,i_{d})$. The total number of unit cells is given by $N_{\mathrm{cell}}:=L_{1}L_{2}\cdots L_{d}$. The Hilbert space of the system is spanned by an orthonormal basis set {|$\{\ket{\vec{i},a}\}$ ($\vec{i}=(i_{1},\cdots,i_{d}),i_{\lambda}=0,\cdots,L_{\lambda}-1$, and $a=1,\cdots,D_{\mathrm{loc}}$), where |$\ket{\vec{i},a}$ denotes a quantum state at site $\vec{i}$ with the index $a$ labeling additional degrees of freedom such as spin, orbit, sublattice, etc. The dimension of the Hilbert space in each unit cell is denoted by $D_{\mathrm{loc}}$, and thus the dimension of the entire Hilbert space $\mathcal{H}$ of the system is given by $D=N_{\mathrm{cell}}D_{\mathrm{loc}}$. (Note that we consider a one-body problem here.) The density matrix of the system can be expanded in terms of the orthonormal basis set as
@@
We impose the periodic boundary condition
@@
Then there is an isomorphic mapping from a CPTP map
@@
Here, we note that the operator $\tilde{c}_{\vec{i},a,c,\vec{\mu}}^{\dagger}$ has two additional indices $c$ and $\vec{\mu}$ compared with the original basis |$\ket{\vec{i},a}$ of the Hilbert space since $\tilde{\mathcal{E}}$ acts on the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$. Physically, this corresponds to the fact that the density matrix has bra and ket degrees of freedom. The non-Hermitian matrix $X(\vec{k})$ is a counterpart of the Bloch Hamiltonian in the band theory, and therefore we shall call it the Bloch matrix. The mode expansion in Eq. (14) of the density matrix after feedback control can be expressed in terms of the eigensystem of the Bloch matrix (see Appendix A.3).
@@ III.2 Topology of quantum channels with local Kraus operators
We are now in a position to discuss topology of feedback control. We consider a homotopy between CPTP maps, that is, two CPTP maps $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$ are homotopic to each other if and only if there exists a continuous one-parameter family $\Xi_{s}$ of CPTP maps (called a homotopy [81]) that preserves the spectral gap and satisfies $\Xi_{0}=\mathcal{E}_{0}$ and $\Xi_{1}=\mathcal{E}_{1}$. Intuitively speaking, the existence of a homotopy between two CPTP maps means that they can continuously be deformed to each other. Here, the spectral gap is either a point gap or a line gap according to the general discussion of topology of non-Hermitian operators [33, 34]. Specifically, a point gap at $\xi=\xi_{\mathrm{PG}}$ means that a CPTP map under consideration does not have an eigenvalue $\xi_{\mathrm{PG}}$. Similarly, a line gap for a line $l_{\mathrm{LG}}$ means that a CPTP map does not have an eigenvalue on $l_{\mathrm{LG}}$. Here we mostly focus on the case of a point gap and discuss the case of a line gap in Sec. VII.5.
@@
In the momentum-space representation in Eq. (28), the eigenspectrum of a CPTP map can be calculated from the Bloch matrix defined in Eq. (31). Thus, a homotopy between translationally invariant CPTP maps is obtained from that between their Bloch matrices. The Bloch matrix $X(\vec{k})$ defines a map from the Brillouin zone (i.e., a $d$-dimensional torus) to a space of non-Hermitian matrices. We can exploit this property to introduce a topological invariant for a CPTP map. If two CPTP maps (or their Bloch matrices) have different values of a topological invariant, they are not homotopic, and cannot continuously be deformed to each other without closing the spectral gap (see Fig. 2). Thus topological invariants can be used to homotopically classify CPTP maps.
@@
Figure 2: Eigenspectra $\{\xi_{n}^{(0)}(\vec{k})\}$ and $\{\xi_{n}^{(1)}(\vec{k})\}$ (shown by blue areas) of two translationally invariant CPTP maps $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$ that can be calculated from their Bloch matrices defined on the Brillouin zone (the $d$-dimensional torus $T^{d}$). If the Bloch matrices have different values of a topological invariant, the two CPTP maps cannot continuously be deformed to each other without closing the spectral gap. In the figure, a point gap at $\xi_{\mathrm{PG}}$ is depicted by a red square.
@@
Here we note that the Bloch matrix is an $N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}\times N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}$ matrix. Since the size of the matrix diverges in the limit of $N_{\mathrm{cell}}\to\infty$, the definition of a topological invariant of a Bloch matrix is nontrivial. Here we show that this problem can be circumvented by imposing locality on the Kraus operators. We assume that the Kraus operators are local in space and have a finite support. Physically, this implies that the interaction between the system and the measurement apparatus (the feedback controller) is local. Let $\ell$ be the diameter of the support of the Kraus operators, and let $R$ be the range of locality of the Kraus operators. Namely, we assume
@@
which indicates that the Bloch matrix $X(\vec{k})$ can be truncated without changing its nonzero eigenvalues and replaced by a matrix $X_{\mathrm{trunc}}(\vec{k})$ with a finite dimension that does not increase with increasing the system size (see Fig. 3). Hence, topology of CPTP maps can be diagnosed by that of the truncated Bloch matrix $X_{\mathrm{trunc}}(\vec{k})$, which remains to be a finite-size matrix in the $N_{\mathrm{cell}}\to\infty$ limit. Topological feedback control is thus characterized by a topological invariant for the truncated Bloch matrix.
Figure 3: Schematic representation of the Bloch matrix $X(\vec{k})$ which is an $N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}\times N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}$ matrix. If the Kraus operators are local and have a finite support, the Bloch matrix can be truncated into a matrix $X_{\mathrm{trunc}}(\vec{k})$ with dimension $D_{\mathrm{trunc}}$ that remains finite in the limit of $N_{\mathrm{cell}}\to\infty$.
An important consequence of the locality of Kraus operators is that the CPTP map has highly degenerate zero eigenvalues. This can be seen by noting that
@@
since the right-hand side cannot have a zero eigenvalue. Thus, the topological classification of quantum channels cannot be reduced to that of Lindbladians discussed in literature [64, 65, 66, 67].
If Kraus operators are not strictly local but have an exponentially decaying tail, topology of a quantum channel may be discussed by approximating them with local ones because topology is insensitive to small deformations. Physically, this should hold true unless rare measurement outcomes with extremely small probabilities due to an exponentially decaying tail of wavefunctions significantly alter the physics. While this is expected to be true for most cases, it would be interesting to find the cases where rare events due to the exponentially small tails of Kraus operators play a significant role in feedback control. We do not further discuss those situations in this paper.
@@ III.3 Stable equivalence of quantum channels
In the definition of the equivalence relation in terms of a homotopy presented in the previous subsection, two CPTP maps acting on systems with different numbers of internal degrees of freedom are not homotopic to each other. A more convenient classification of topological phases is obtained by introducing the notion of stable equivalence [18, 27]. In the topological band theory, two band insulators are stably equivalent if they are continuously connected to each other after an addition of some trivial bands [18]. Analogously, we define the stable equivalence of CPTP maps as follows. First, we generalize the quantum states at each unit cell as |$\ket{\vec{j},a}$ ($a=1,\cdots,D_{\mathrm{loc}},D_{\mathrm{loc}}+1,\cdots,D_{\mathrm{loc}}+n_{1}$) by adding $n_{1}$ extra local degrees of freedom, which can physically be interpreted as ancillary sublattice sites. We assume that the Kraus operators $\hat{K}_{m}$ act trivially on the extra degrees of freedom, i.e., $\hat{K}_{m}\ket{\vec{j},a}=0$ for $a=D_{\mathrm{loc}}+1,\cdots,D_{\mathrm{loc}}+n_{1}$. Then, we introduce a projective measurement of the ancillary degrees of freedom, which is described by
$\mathcal{P}_{n_{1}}(\hat{\rho}):=\sum_{\vec{j}}\sum_{a=D_{\mathrm{loc}}+1}^{n_{1}}\hat{P}_{\vec{j},a}\hat{\rho}\hat{P}_{\vec{j},a},$ (40)
where $\hat{P}_{\vec{j},a}:=\ket{\vec{j},a}\bra{\vec{j},a}$ | is the projection operator onto the $a$th internal degree of freedom at site $\vec{j}$. The projective measurement channel (40) plays a role of “trivial bands” in the sense that all of its nonzero eigenvalues are unity (see Appendix C). We note that the projection operators $\hat{P}_{\vec{j},a}$ are local in space in accordance with the discussion in Sec. III.2. We define the direct sum of a quantum channel $\mathcal{E}(\hat{\rho})=\sum_{m}\hat{K}_{m}\hat{\rho}\hat{K}_{m}^{\dagger}$ for the system and the projective measurement channel (40) for the extra bands by
$(\mathcal{E}\oplus\mathcal{P}_{n_{1}})(\hat{\rho}):=\sum_{m}\hat{K}_{m}\hat{\rho}\hat{K}_{m}^{\dagger}+\sum_{\vec{j}}\sum_{a=D_{\mathrm{loc}}+1}^{n_{1}}\hat{P}_{\vec{j},a}\hat{\rho}\hat{P}_{\vec{j},a}.$ (41)
Let $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ be CPTP maps. Then, the CPTP maps $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ are said to be stably equivalent if and only if $\mathcal{E}_{1}\oplus\mathcal{P}_{n_{1}}$ are homotopic to $\mathcal{E}_{2}\oplus\mathcal{P}_{n_{2}}$ for some two non-negative integers $n_{1}$ and $n_{2}$ (see Fig. 4).
Figure 4: Stable equivalence of CPTP maps. A CPTP map with the eigenspectrum shown by blue areas in the top-left figure is not homotopic to the other CPTP map with the eigenspectrum shown by blue and green areas in the bottom figure, since the numbers of bands (eigenvalues) are different. However, they can be made homotopic if we add a trivial CPTP map with the eigenspectrum shown in the top-right figure so that the numbers of eigenvalues become equal to each other.
@@ IV Topological feedback control
IV.1 Chiral Maxwell’s demon
Having established the general framework of topology of quantum channels, here we present a prototypical example of topological quantum feedback control characterized by a topological invariant of the Bloch matrix of a quantum channel. Our model is a quantum counterpart of Maxwell’s demon in Ref. [54] (see also Ref. [82]). We consider a particle hopping on a one-dimensional (1D) lattice with $L$ sites, where a quantum state at site $i$ is denoted by |$\ket{i}$. We first perform a projective measurement of the position of this particle with measurement operators $\hat{M}_{m}=\ket{m}\bra{m}$ | $(m=1,\cdots,L)$. If the measurement outcome is $j$, we raise the potential at site $j-1$ and perform the feedback control with the Hamiltonian
@@ IV.1 Chiral Maxwell’s demon
where $J\in\mathbb{R}$ is the hopping amplitude and $V>0$ is the height of the potential. The unitary operator for the feedback control is given by $\hat{U}_{j}=\exp[-i\hat{H}_{j}\tau]$, where $\tau$ is the duration of the feedback control. The potential barrier placed at site $j-1$ prevents the particle from going to the left, thereby achieving unidirectional (chiral) transport of the particle to the right (see Fig. 5). Here, we do not need any external field to induce the transport in the system; instead, the information gained from measurement is used to drive the particle unidirectionally. The feedback control thus rectifies quantum fluctuations in measurement outcomes and convert them into a unidirectional transport in a manner similar to an information ratchet [83]. We therefore refer to this model as the chiral Maxwell’s demon.
Figure 5: Schematic illustration of the chiral Maxwell’s demon. A quantum-mechanical particle hops on a one-dimensional lattice with hopping amplitude $J$. We perform a projective measurement of the position of the particle. If we find the particle at site $j$, we raise the potential at site $j-1$ to suppress the probability of the particle hopping to the left.
The eigenvalues $\xi_{n}(k)\ (n=0,\cdots,L-1)$ of a quantum channel of this feedback control can be calculated by using the method described in Appendix B and are given by
@@
where $\xi_{\mathrm{PG}}$ denotes the location of a point gap [33]. Here, the eigenvalues $\xi_{n}(k)=0\ (n\neq 0)$ do not contribute to the winding number (since $\partial_{k}\xi_{n}(k)=0$) and thus are excluded from Eq. (46). The winding number takes an integer value and it is invariant under continuous deformation of a feedback control protocol as long as the point gap at $\xi_{\mathrm{PG}}$ is not closed. This is consistent with the $\mathbb{Z}$ classification of non-Hermitian matrices in 1D systems according to stable equivalence [33].

### v3

@@ III Topological equivalence of quantum channels
In this section, we introduce the topology of quantum channels (CPTP maps) to investigate topological phases induced by feedback control.
@@ III.1 Setup
be a set of lattice sites, where $\vec{a}_{\lambda}\ (\lambda=1,\cdots,d)$ denote the primitive lattice vectors. Each lattice site $\vec{R}_{\vec{i}}$ is specified by a set of integers $\vec{i}=(i_{1},\cdots,i_{d})$. The total number of unit cells is given by $N_{\mathrm{cell}}:=L_{1}L_{2}\cdots L_{d}$. The Hilbert space of the system is spanned by an orthonormal basis set {|$\{\ket{\vec{i},a}\}$ ($\vec{i}=(i_{1},\cdots,i_{d}),i_{\lambda}=0,\cdots,L_{\lambda}-1$, and $a=1,\cdots,D_{\mathrm{loc}}$), where |$\ket{\vec{i},a}$ denotes a quantum state at site $\vec{i}$, with the index $a$ labeling additional degrees of freedom such as spin, orbit, sublattice, etc. The dimension of the Hilbert space in each unit cell is denoted by $D_{\mathrm{loc}}$, and thus the dimension of the entire Hilbert space $\mathcal{H}$ of the system is given by $D=N_{\mathrm{cell}}D_{\mathrm{loc}}$. (Note that we consider a one-body problem here.) The density matrix of the system can be expanded in terms of the orthonormal basis set as
@@
We impose the periodic boundary condition (PBC)
@@
Correspondingly, there is an isomorphic mapping from a CPTP map
@@
Here, we note that the operator $\tilde{c}_{\vec{i},a,c,\vec{\mu}}^{\dagger}$ has two additional indices $c$ and $\vec{\mu}$ compared with the original basis |$\ket{\vec{i},a}$ of the Hilbert space since $\tilde{\mathcal{E}}$ acts on the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$. Physically, this point corresponds to the fact that the density matrix has bra and ket degrees of freedom. The non-Hermitian matrix $X(\vec{k})$ is a counterpart of the Bloch Hamiltonian in the band theory, and therefore we call it the Bloch matrix. The mode expansion in Eq. (14) of the density matrix after feedback control can be expressed in terms of the eigensystem of the Bloch matrix (see Appendix A.3).
@@ III.2 Topology of quantum channels with local Kraus operators
We are now in a position to discuss the topology of feedback control. We consider a homotopy between CPTP maps, that is, two CPTP maps $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$ are homotopic to each other if and only if there exists a continuous one-parameter family $\Xi_{s}$ of CPTP maps (called a homotopy [81]) that preserves the spectral gap and satisfies $\Xi_{0}=\mathcal{E}_{0}$ and $\Xi_{1}=\mathcal{E}_{1}$. Intuitively speaking, the existence of a homotopy between two CPTP maps means that they can continuously be deformed into each other. Here, the spectral gap is either a point gap or a line gap according to the general discussion of topology of non-Hermitian operators [33, 34]. Specifically, a point gap at $\xi=\xi_{\mathrm{PG}}$ means that the CPTP map under consideration does not have an eigenvalue $\xi_{\mathrm{PG}}$. Similarly, a line gap for a line $l_{\mathrm{LG}}$ means that a CPTP map does not have an eigenvalue on $l_{\mathrm{LG}}$. Here, we mostly focus on the case of a point gap and discuss the case of a line gap in Sec. VII.5.
@@
In the momentum-space representation in Eq. (28), the eigenspectrum of a CPTP map can be calculated from the Bloch matrix defined in Eq. (31). Thus, a homotopy between translationally invariant CPTP maps is obtained from that between their Bloch matrices. The Bloch matrix $X(\vec{k})$ defines a map from the Brillouin zone (i.e., a $d$-dimensional torus) to a space of non-Hermitian matrices. We can exploit this property to introduce a topological invariant for a CPTP map. If two CPTP maps (or their Bloch matrices) have different values of a topological invariant, they are not homotopic and cannot continuously be deformed into each other without closing the spectral gap (see Fig. 2). Thus, topological invariants can be used to homotopically classify CPTP maps.
@@
Figure 2: Eigenspectra $\{\xi_{n}^{(0)}(\vec{k})\}$ and $\{\xi_{n}^{(1)}(\vec{k})\}$ (shown by blue areas) of two translationally invariant CPTP maps $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$ that can be calculated from their Bloch matrices defined on the Brillouin zone (the $d$-dimensional torus $T^{d}$). If the Bloch matrices have different values of a topological invariant, the two CPTP maps cannot continuously be deformed into each other without closing the spectral gap. In the figure, a point gap at $\xi_{\mathrm{PG}}$ is depicted by a red square.
@@
Here, we note that the Bloch matrix is an $N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}\times N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}$ matrix. Since the size of the matrix diverges in the limit $N_{\mathrm{cell}}\to\infty$, the definition of a topological invariant of a Bloch matrix is nontrivial. We show that this problem can be circumvented by imposing locality on the Kraus operators. We assume that the Kraus operators are local in space and have a finite support. Physically, this assumption implies that the interaction between the system and the measurement apparatus (the feedback controller) is local. Let $\ell$ be the diameter of the support of the Kraus operators, and let $R$ be the range of locality of the Kraus operators. Specifically, we assume
@@
which indicates that the Bloch matrix $X(\vec{k})$ can be truncated without changing its nonzero eigenvalues and can be replaced by a matrix $X_{\mathrm{trunc}}(\vec{k})$ with a finite dimension that does not increase when increasing the system size (see Fig. 3). Hence, the topology of a CPTP map can be diagnosed by that of its truncated Bloch matrix $X_{\mathrm{trunc}}(\vec{k})$, which remains a finite-size matrix in the $N_{\mathrm{cell}}\to\infty$ limit. Topological feedback control is thus characterized by a topological invariant for the truncated Bloch matrix.
Figure 3: Schematic representation of the Bloch matrix $X(\vec{k})$ which is an $N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}\times N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}$ matrix. If the Kraus operators are local and have a finite support, the Bloch matrix can be truncated into a matrix $X_{\mathrm{trunc}}(\vec{k})$ with dimension $D_{\mathrm{trunc}}$ that remains finite in the limit $N_{\mathrm{cell}}\to\infty$.
An important consequence of the locality of Kraus operators is that the CPTP map has highly degenerate zero eigenvalues. This result can be seen by noting that
@@
since the right-hand side cannot have a zero eigenvalue. Thus, the topological classification of quantum channels cannot be reduced to that of the Lindbladians discussed in the literature [64, 65, 66, 67].
If Kraus operators are not strictly local but have an exponentially decaying tail, the topology of a quantum channel may be discussed by approximating them with local ones because topology is insensitive to small deformations. Physically, this case should hold true unless rare measurement outcomes with extremely small probabilities due to an exponentially decaying tail of wavefunctions significantly alter the physics. While it is expected to be true for most cases, it would be interesting to find the cases where rare events due to the exponentially small tails of Kraus operators play a significant role in feedback control. We do not discuss those situations further in this paper.
@@ III.3 Stable equivalence of quantum channels
In the definition of the equivalence relation in terms of a homotopy presented in the previous subsection, two CPTP maps acting on systems with different numbers of internal degrees of freedom are not homotopic to each other. A more convenient classification of topological phases is obtained by introducing the notion of stable equivalence [18, 27]. In the topological band theory, two band insulators are stably equivalent if they are continuously connected to each other after the addition of some trivial bands [18]. Analogously, we define the stable equivalence of CPTP maps as follows. First, we generalize the quantum states at each unit cell as |$\ket{\vec{j},a}$ ($a=1,\cdots,D_{\mathrm{loc}},D_{\mathrm{loc}}+1,\cdots,D_{\mathrm{loc}}+n_{1}$) by adding $n_{1}$ extra local degrees of freedom, which can physically be interpreted as ancillary sublattice sites. We assume that the Kraus operators $\hat{K}_{m}$ act trivially on the extra degrees of freedom, i.e., $\hat{K}_{m}\ket{\vec{j},a}=0$ for $a=D_{\mathrm{loc}}+1,\cdots,D_{\mathrm{loc}}+n_{1}$. Next, we introduce a projective measurement of the ancillary degrees of freedom, which is described by
$\mathcal{P}_{n_{1}}(\hat{\rho}):=\sum_{\vec{j}}\sum_{a=D_{\mathrm{loc}}+1}^{D_{\mathrm{loc}}+n_{1}}\hat{P}_{\vec{j},a}\hat{\rho}\hat{P}_{\vec{j},a},$ (40)
where $\hat{P}_{\vec{j},a}:=\ket{\vec{j},a}\bra{\vec{j},a}$ | is the projection operator onto the $a$th internal degree of freedom at site $\vec{j}$. The projective measurement channel (40) plays the role of “trivial bands” in the sense that all of its nonzero eigenvalues are unity (see Appendix C). We note that the projection operators $\hat{P}_{\vec{j},a}$ are local in space in accordance with the discussion in Sec. III.2. We define the direct sum of a quantum channel $\mathcal{E}(\hat{\rho})=\sum_{m}\hat{K}_{m}\hat{\rho}\hat{K}_{m}^{\dagger}$ for the system and the projective measurement channel (40) for the extra bands by
$(\mathcal{E}\oplus\mathcal{P}_{n_{1}})(\hat{\rho}):=\sum_{m}\hat{K}_{m}\hat{\rho}\hat{K}_{m}^{\dagger}+\sum_{\vec{j}}\sum_{a=D_{\mathrm{loc}}+1}^{D_{\mathrm{loc}}+n_{1}}\hat{P}_{\vec{j},a}\hat{\rho}\hat{P}_{\vec{j},a}.$ (41)
Let $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ be CPTP maps. Then, the CPTP maps $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ are said to be stably equivalent if and only if $\mathcal{E}_{1}\oplus\mathcal{P}_{n_{1}}$ are homotopic to $\mathcal{E}_{2}\oplus\mathcal{P}_{n_{2}}$ for two non-negative integers $n_{1}$ and $n_{2}$ (see Fig. 4).
Figure 4: Stable equivalence of CPTP maps. A CPTP map with the eigenspectrum shown by blue areas in the top-left figure is not homotopic to the other CPTP map with the eigenspectrum shown by blue and green areas in the bottom figure since the numbers of bands (eigenvalues) are different. However, they can be made homotopic if we add a trivial CPTP map with the eigenspectrum shown in the top-right figure so that the numbers of eigenvalues become equal to each other.
@@ IV Topological feedback control
IV.1 Chiral Maxwell demon
Having established the general framework of topology of quantum channels, here we present a prototypical example of topological quantum feedback control characterized by a topological invariant of the Bloch matrix of a quantum channel. Our model is a quantum counterpart of Maxwell’s demon in Ref. [54] (see also Ref. [82]). We consider a particle hopping on a one-dimensional (1D) lattice with $L$ sites, where a quantum state at site $i$ is denoted by |$\ket{i}$. We first perform a projective measurement of the position of this particle with measurement operators $\hat{M}_{m}=\ket{m}\bra{m}$ | $(m=1,\cdots,L)$. If the measurement outcome is $j$, we raise the potential at site $j-1$ and perform feedback control with the Hamiltonian
@@ IV.1 Chiral Maxwell demon
where $J\in\mathbb{R}$ is the hopping amplitude and $V>0$ is the height of the potential. The unitary operator for feedback control is given by $\hat{U}_{j}=\exp[-i\hat{H}_{j}\tau]$, where $\tau$ is the duration of feedback control. The potential barrier placed at site $j-1$ prevents the particle from going to the left, thereby achieving unidirectional (chiral) transport of the particle to the right (see Fig. 5). Here, we do not need any external field to cause the transport in the system; instead, the information gained from the measurement is used to drive the particle unidirectionally. Feedback control thus rectifies quantum fluctuations in measurement outcomes and converts them into a unidirectional transport in a manner similar to an information ratchet [83]. We therefore refer to this model as the chiral Maxwell demon.
Figure 5: Schematic illustration of the chiral Maxwell demon. A quantum-mechanical particle hops on a one-dimensional lattice with hopping amplitude $J$. We perform a projective measurement of the position of the particle. If we find the particle at site $j$, we raise the potential at site $j-1$ to suppress the probability of the particle hopping to the left.
The eigenvalues $\xi_{n}(k)\ (n=0,\cdots,L-1)$ of a quantum channel of this feedback control can be calculated by using the method described in Appendix B. They are given by
@@
where $\xi_{\mathrm{PG}}$ denotes the location of a point gap [33]. Here, the eigenvalues $\xi_{n}(k)=0\ (n\neq 0)$ do not contribute to the winding number [since $\partial_{k}\xi_{n}(k)=0$] and thus are excluded from Eq. (46). The winding number takes an integer value and it is invariant under continuous deformation of a feedback control protocol as long as the point gap at $\xi_{\mathrm{PG}}$ is not closed. This case is consistent with the $\mathbb{Z}$ classification of non-Hermitian matrices in 1D systems according to stable equivalence [33].